# Directed Hypercube Routing, a Generalized Lehman-Ron Theorem, and Monotonicity Testing

### v1

@@ 1 Introduction
Two subsets $S,T$ of $V(H)$ are called a matched pair if there exists a bijection $\phi:S\to T$ such that $s\prec\phi(s)$ for all $s\in S$; we denote a matched pair by $(S,T;\phi)$. An early writeup of Goldreich-Goldwasser-Ron [GGR] posed a routing question, inspired by questions in monotonicity testing, which was solved by Lehman and Ron [LR01]. (More discussion in §4.) They were interested in the following natural question: given any matched pair $(S,T;\phi)$, can one find (edge or vertex) disjoint directed paths111 In their paper, Lehman and Ron consider these paths to be disjoint chains of subsets, as reflected in their paper title. from $S$ to $T$? Remarkably, they proved that if all points in $S$ (resp. $T$) have the same Hamming weight, then the answer is affirmative: one can find vertex disjoint paths from $S$ to $T$! More precisely, for an integer $0\leq i\leq d$, let $L_{i}$ denote the $i$th layer of $H$, that is, $L_{i}:=\{x\in V(H)~{}:~{}||x||_{1}=i\}$. We refer to this beautiful statement as the “Lehman-Ron (LR) theorem”.
@@
Lehman-Ron’s proof of Theorem 1.1 is by induction on $|S|$ and on the quantity $(j-i)$, the distance between the layers in which $S$ and $T$ lie. The base case of $j-i=1$ is obvious as the bijection gives us the matching between $S$ and $T$. The heart of the proof essentially shows the existence of a set of vertices $U$ either in layer $L_{j-1}$ or $L_{i+1}$ and two bijections $\phi^{\prime}:S\to U$ and $\phi^{\prime\prime}:U\to T$ such that $(S,U;\phi^{\prime})$ and $(U,T;\phi^{\prime\prime})$ are matched pairs. This last part is a neat argument which uses Menger’s theorem, which is a special case of the max-flow-min-cut theorem, on an auxiliary graph that they create. But how can one get two edge disjoint LR solutions? The reader may notice that even the “base case” of $j-i=2$, that is, when $S$ and $T$ are two levels apart is itself non-trivial (indeed, we don’t really know a much simpler way to solve this than the general case). And so, a new idea is needed to prove Theorem 1.2.
@@ 2 Alternate proof of the Lehman-Ron Theorem
In other words, the cover graph is the subset of the hypercube, that contains all vertices $v$ such that $s\prec v\prec t$ (for $s\in S,t\in T$). The cover graph inherits “layers” via intersection with the original hypercube layers. In particular, layer $L_{i}$ of the cover graph is only $S$ and layer $L_{j}$ is only $T$.
@@
There exists a collection of vertex disjoint paths $\mathcal{P}$ where every path $p\in\mathcal{P}$ begins at a vertex in $S$ and ends at a vertex in $p$ and
@@
Using the above collection of paths, we make a key definition.
@@
Now, let’s see how to get a gateway in the next layer $L_{i+1}$. Consider any edge $(s,x^{{(1)}})$ in $G_{S,T}$, and suppose this edge corresponds to projecting according to some dimension $r$. That is $s_{r}=0$ and $x^{(1)}_{r}=1$. If $x^{(1)}$ is not in any path in $\mathcal{P}$, we have discovered the desired gateway in $L_{i+1}$. Otherwise, $x^{(1)}$ lies on some path, say, $P\in\mathcal{P}$. Follow $P$ forwards for a single edge from $x^{(1)}$, to get to $x^{(2)}$. Note that $x^{(2)}\in L_{i+2}$ and has $r$th-coordinate $1$. Now, one can project “down” on the $r$-coordinate to get $x^{(3)}\in L_{i+1}$. Observe that $(s,x^{(3)})$ is a projection of the edge $(x^{(1)},x^{(2)})$ along the $r$th-dimension and hence is an edge of the cover-graph. Next, observe that $x^{(3)}$ cannot be in ${\cal T}$, so either $x^{(3)}\in\mathcal{S}$ or $x^{(3)}\in C$. If $x_{3}\in\mathcal{S}$ and not on any path in $\mathcal{P}$, we are done. Otherwise $x^{(3)}$ lies on some other path $Q\in\mathcal{P}$. We now walk backwards along $Q$, to get $x^{(4)}\in S$. Noting that $x^{(4)}$ has $r$-coordinate 00, we can redo the entire process above. Observe that each “step” proceeds along a matching. Either we project, walk from $L_{i+1}$ to $L_{i+2}$ using a path in $\mathcal{P}$, or walk backward from $L_{i+1}$ to $L_{i}$ using a path in $\mathcal{P}$. Each of these is using a matching edge, and no vertex is ever visited twice in the entire process. Hence, this process must terminate, at which point a gateway is discovered. And then one uses the same idea to obtain a gateway vertex in $L_{i+2}$, and so on. One can convert this idea into a formal proof, but it becomes notationally cumbersome. A cleaner proof method is to consider a potential “fixed point” of this process, and prove a contradiction.
@@
As argued above, $L_{i}$ has a gateway vertex. Let $k\in[i,j-1]$ be the largest value such that $L_{k}$ contains a gateway vertex. If $k=j-1$, we are done. So suppose that $k<j-1$. We now engineer a contradiction. Let $v^{\star}$ be a gateway in $L_{k}$. So $v^{\star}\in\mathcal{S}$ and has an edge $(v^{\star},w)$ leaving it. Let $r$ be the dimension of this edge implying $v^{\star}_{r}=0$ and $w_{r}=1$. Let $\Pi_{r}$ be the projection operator which flips the $r$th coordinate; so $\Pi_{r}(v^{\star})=w$ and vice versa. Define the following sets; we give a illustration for convenience where the pink highlighted edges participate in paths of $\mathcal{P}$.
@@
  • $B=\{b~{}:~{}(x,b)\in P,~{}\text{for some}~{}P\in\mathcal{P}\}$.
@@
Observe that $X\cup Y\in L_{k+1}$, and $B\in L_{k+2}$. We next argue these subsets lie in the cover-graph. By definition of $A$, $X\subseteq G_{\mathcal{S},{\cal T}}$, and since $B$ is obtained via paths from $X$, we get $B\subseteq G_{\mathcal{S},{\cal T}}$. Finally, for any $y\in Y$, note that $b=\Pi_{r}(y)$ lies in $B$, and there’s a path $(a,x,b)$ for some $a\in A$ and $x=\Pi_{r}(a)$. But note that $(a,y)$ edge is the $r$-projection of $(x,b)$, and so $(a,y,b)$ is a path implying $y\in G_{\mathcal{S},{\cal T}}$. Thus, $Y$ also lies in the cover-graph.
@@
Next consider any $y\in Y$. As argued above, there is a vertex $a\in A$ such that $(a,y)$ is an edge in the cover-graph (the projection of $(x,b)\in P$ where $b=\Pi_{r}(y)$). Since $y\in L_{k+1}$, by our choice of $k$, $y$ isn’t a gateway vertex. Since $(a,y)$ is an edge, we must have $y\in\mathcal{S}$ or $y\in C$. In the former case since $y$ isn’t a gateway vertex, and in the latter case by Lemma 2.3, there exists $Q\in\mathcal{P}$ with $y\in Q$. Let $(a^{\prime},y)$ be the edge in $Q$ taking $y$ “backward” along $Q$. We claim that $a^{\prime}\in A$. If $y\in\mathcal{S}$, then $a^{\prime}\in\mathcal{S}$ since $(a^{\prime},y)$ is an edge; if $y\in C$ then by Lemma 2.3, $a^{\prime}\in\mathcal{S}$ (the path $Q$ doesn’t contain two cut-vertices). Since $(a^{\prime},x^{\prime}=\Pi_{r}(a^{\prime}))$ lies in the cover graph since the $r$-projection of the $(a^{\prime},y)$ lies there, we get $x^{\prime}$ lies in the cover-graph, implying $a^{\prime}$ must lie in $A$. The above figure illustrates this. In short, there are $|Y|$ paths of $\mathcal{P}$ that contain vertices of $A$. Since $|Y|=|A|$ and since these paths are all vertex-disjoint, we conclude all vertices in $A$ are present in some path of $\mathcal{P}$. However, the gateway vertex $v^{\star}\in A$ and doesn’t lie on any path of $\mathcal{P}$. Contradiction. ∎
@@
  • $B=\{b~{}:~{}(x,b)\in P,~{}\text{for some}~{}P\in\mathcal{P}\}$.
@@
As in the proof of Lemma 2.5, we have $X\cup Y\in L_{k+1}$, and $B\in L_{k+2}$ and that $A,X,B,Y$ all lie in the cover graph $G_{\mathcal{S},{\cal T}}$. As there, we crucially use the property that the projection of every $(A,Y)$ edge lies is an $(X,B)$ edge. Note that, we have $|A|=|X|$ and $|B|=|Y|$. However, unlike in Lemma 2.5, there can be two paths of $\mathcal{P}$ that may contain $x\in X$, and thus, $|B|$ may be larger in size than $|X|$. And this makes the proof slightly more complicated.
We now restrict our attention to these four sets. As a result, we only consider the cover graph $G_{A,B}$, which, recall, is the union of all paths from vertices in $A$ to vertices in $B$ (irrespective of whether they contain cut-elements). Both $A$ and $B$ lie in this graph by definition. We assert $X\subseteq G_{A,B}$. Take a vertex $x\in X$. If $x\in P$ for some path $P\in\mathcal{P}$ then it lies in $G_{A,B}$ by construction. We now show if $x$ is not any path in $\mathcal{P}$, we already have a contradiction. Indeed, the edge $(a,x)$ where $a=\Pi_{r}(x)$ isn’t in $F$ for otherwise, by Lemma 3.2 there would be a path containing $(a,x)$. Similarly $x\notin C$. So, $x\in S$. Since $x\in L_{k+1}$ and $k<j-1$, there is at least one edge leaving it, and as before, this edge isn’t in $F$. So, $x$ is a gateway vertex which contradicts the choice of $k$. In sum, the subset $X$ lies in $G_{A,B}$. And since every $(A,Y)$ edge is an $r$-projection of an $(X,B)$ edge, we argue that $Y$ also lies in $G_{A,B}$.
@@ 4 Connections to monotonicity testing
The motivation for Conjecture 1.3 (and indeed, Theorem 1.1) is a deeper understanding of the problem of monotonicity testing of functions, a problem which, especially over the hypercube and hypergrid domains, has had a rich history of more than 25 years  [Ras99, EKK+00, GGL+00, DGL+99, LR01, FLN+02, HK03, AC06, HK08, ACCL07, Fis04, SS08, Bha08, BCSM12, FR10, BBM12, RRS+12, BGJ+12, CS13, CS14, CST14, BRY14a, BRY14b, CDST15, CDJS17, KMS18, BB21, CWX17, BCS18, BCS20, BKR20, HY22, BKKM23, BCS23]. A function $f:\{0,1\}^{d}\to\{0,1\}$ is monotone if $\forall x,y\in\{0,1\}^{d}$ where $x\prec y$, $f(x)\leq f(y)$. The distance between two functions $f,g$ is |$|\{x:f(x)\neq g(x)\}|/2^{d}$, and the distance of $f$ to monotonicity, denote $\varepsilon_{f}$, is the minimum distance of $f$ to a monotone $g$. A function $f$ is said to be $\varepsilon$-far from monotone if $\varepsilon_{f}\geq\varepsilon$. The aim of a tester is to distinguish a monotone function from one that is “far” from monotone. There is a special focus on non-adaptive monotonicity testers with one-sided error. These are testers that (i) always accept monotone functions, and (ii) make all their queries in advance. After a long line of work, this has been resolved (up to $\log d,\mathrm{poly}(\varepsilon^{-1})$ factors) to be $\Theta(\sqrt{d})$ [GGL+00, FLN+02, CS14, CST14, KMS18, CDST15, CWX17].
@@
In the next subsection, we give combinatorial interpretations to each of these statements. The reason for Conjecture 1.3 and a deeper study of hypercube routing was to get alternate proofs of Theorem 4.4. A big mystery of all these directed isoperimetric inequalities is the appearance of $\varepsilon_{f}$, the distance to monotonicity, as a “directed version” of the variance of $f$. It appears as if $\varepsilon_{f}$ is the “right measure” of directed volume. We hope that alternate proofs of Theorem 4.4 may shed some light on this mystery.

### v2

@@ 1 Introduction
Two subsets $S,T$ of $V(H)$ are called a matched pair if there exists a bijection $\phi:S\to T$ such that $s\prec\phi(s)$ for all $s\in S$; we denote a matched pair by $(S,T;\phi)$. An early writeup of Goldreich-Goldwasser-Ron [GGR] posed a routing question, inspired by questions in monotonicity testing, which was solved by Lehman and Ron [LR01]. (More discussion in §4.) They were interested in the following natural question: given any matched pair $(S,T;\phi)$, can one find (edge or vertex) disjoint directed paths111 In their paper, Lehman and Ron consider these paths to be disjoint chains of subsets. from $S$ to $T$? Remarkably, they proved that if all points in $S$ (resp. $T$) have the same Hamming weight, then the answer is affirmative: one can find vertex disjoint paths from $S$ to $T$! More precisely, for an integer $0\leq i\leq d$, let $L_{i}$ denote the $i$th layer of $H$, that is, $L_{i}:=\{x\in V(H)~{}:~{}||x||_{1}=i\}$. We refer to this beautiful statement as the ”Lehman-Ron (LR) theorem”.
@@
Lehman-Ron’s proof of Theorem 1.1 is by induction on $|S|$ and on the quantity $(j-i)$, the distance between the layers in which $S$ and $T$ lie. The base case of $j-i=1$ is obvious as the bijection gives us the matching between $S$ and $T$. The heart of the proof essentially shows the existence of a set of vertices $U$ either in layer $L_{j-1}$ or $L_{i+1}$ and two bijections $\phi^{\prime}:S\to U$ and $\phi^{\prime\prime}:U\to T$ such that $(S,U;\phi^{\prime})$ and $(U,T;\phi^{\prime\prime})$ are matched pairs. This last part is a neat argument which uses Menger’s theorem, which is a special case of the max-flow-min-cut theorem, on an auxiliary graph that they create. But how can one get two edge disjoint LR solutions? The reader may notice that even the ”base case” of $j-i=2$, that is, when $S$ and $T$ are two levels apart is itself non-trivial (indeed, we don’t really know a much simpler way to solve this than the general case). And so, a new idea is needed to prove Theorem 1.2.
@@
Other LR connections. Recent work has generalized the LR theorem different directions in [BGJJ24]. These results find vertex disjoint paths that ”cover” any collection of points specifying certain properties. Consider a subset $X$ of the hypercube that is partitioned into subsets of paths (or chains). Meaning, we partition $X=\bigcup_{i}X_{i}$ such that, the vertices of $X_{i}$ can be ordered according to $\prec$. (Moreover, this is the partition that minimizes the number of sets.) In the vanilla LR setting, each $X_{i}$ is just $(s,\phi(s))$ for each $s\in S$. The main theorem of [BGJJ24] shows that $X$ can be covered by a collection of vertex disjoint paths. A nice implication of their result is that Theorem 1.1 holds even if $S$ and $T$ were not contained in levels, but were antichains.
We also note that the routing perspective in §4 answers a question of Sachdeva from a collection of open problems on Boolean functions [FHH+14] (Pg 19, ”Routing on the hypercube”). We discuss more in §4.
@@ 2 Alternate proof of the Lehman-Ron Theorem
In other words, the cover graph is the subset of the hypercube, that contains all vertices $v$ such that $s\prec v\prec t$ (for $s\in S,t\in T$). The cover graph inherits ”layers” via intersection with the original hypercube layers. In particular, layer $L_{i}$ of the cover graph is only $S$ and layer $L_{j}$ is only $T$.
@@
There exists a collection of vertex disjoint paths $\mathcal{P}$ where every path $p\in\mathcal{P}$ begins at a vertex in $S$ and ends at a vertex in $T$ and
@@
Note that all these paths are in the cover graph $G_{S,T}$. Using the above collection of paths, we make a key definition.
@@
Now, let’s see how to get a gateway in the next layer $L_{i+1}$. Consider any edge $(s,x^{{(1)}})$ in $G_{S,T}$, and suppose this edge corresponds to projecting according to some dimension $r$. That is $s_{r}=0$ and $x^{(1)}_{r}=1$. If $x^{(1)}$ is not in any path in $\mathcal{P}$, we have discovered the desired gateway in $L_{i+1}$. Otherwise, $x^{(1)}$ lies on some path, say, $P\in\mathcal{P}$. Follow $P$ forwards for a single edge from $x^{(1)}$, to get to $x^{(2)}$. Note that $x^{(2)}\in L_{i+2}$ and has $r$th-coordinate $1$. Now, one can project ”down” on the $r$-coordinate to get $x^{(3)}\in L_{i+1}$. Observe that $(s,x^{(3)})$ is a projection of the edge $(x^{(1)},x^{(2)})$ along the $r$th-dimension and hence is an edge of the cover-graph. Next, observe that $x^{(3)}$ cannot be in ${\cal T}$, so either $x^{(3)}\in\mathcal{S}$ or $x^{(3)}\in C$. If $x_{3}\in\mathcal{S}$ and not on any path in $\mathcal{P}$, we are done. Otherwise $x^{(3)}$ lies on some other path $Q\in\mathcal{P}$. We now walk backwards along $Q$, to get $x^{(4)}\in S$. Noting that $x^{(4)}$ has $r$-coordinate 00, we can redo the entire process above. Observe that each ”step” proceeds along a matching. Either we project, walk from $L_{i+1}$ to $L_{i+2}$ using a path in $\mathcal{P}$, or walk backward from $L_{i+1}$ to $L_{i}$ using a path in $\mathcal{P}$. Each of these is using a matching edge, and no vertex is ever visited twice in the entire process. Hence, this process must terminate, at which point a gateway is discovered. And then one uses the same idea to obtain a gateway vertex in $L_{i+2}$, and so on. One can convert this idea into a formal proof, but it becomes notationally cumbersome. A cleaner proof method is to consider a potential ”fixed point” of this process, and prove a contradiction.
@@
Fix a collection of paths $\mathcal{P}$ as given by Lemma 2.3. As argued above, $L_{i}$ has a gateway vertex. Let $k\in[i,j-1]$ be the largest value such that $L_{k}$ contains a gateway vertex. If $k=j-1$, we are done. So suppose that $k<j-1$. We now engineer a contradiction. Let $v^{\star}$ be a gateway in $L_{k}$. So $v^{\star}\in\mathcal{S}$ and has an edge $(v^{\star},w)$ leaving it. Let $r$ be the dimension of this edge implying $v^{\star}_{r}=0$ and $w_{r}=1$. Let $\Pi_{r}$ be the projection operator which flips the $r$th coordinate; so $\Pi_{r}(v^{\star})=w$ and vice versa. Define the following sets; we give a illustration for convenience where the pink highlighted edges participate in paths of $\mathcal{P}$.
@@
  • $B=\{b~{}:~{}(x,b)\in P,~{}\text{for some}~{}x\in X,P\in\mathcal{P}\}$.
@@
Observe that $X\cup Y\subseteq L_{k+1}$, and $B\subseteq L_{k+2}$. We next argue these subsets lie in the cover-graph. By definition of $A$, $X\subseteq G_{\mathcal{S},{\cal T}}$, and since $B$ is obtained via paths from $X$, we get $B\subseteq G_{\mathcal{S},{\cal T}}$. Finally, for any $y\in Y$, note that $b=\Pi_{r}(y)$ lies in $B$, and there’s a path $(a,x,b)$ for some $a\in A$ and $x=\Pi_{r}(a)$. But note that $(a,y)$ edge is the $r$-projection of $(x,b)$, and so $(a,y,b)$ is a path implying $y\in G_{\mathcal{S},{\cal T}}$. Thus, $Y$ also lies in the cover-graph.
@@
Next consider any $y\in Y$. As argued above, there is a vertex $a\in A$ such that $(a,y)$ is an edge in the cover-graph (the projection of $(x,b)\in P$ where $b=\Pi_{r}(y)$). Since $y\in L_{k+1}$, by our choice of $k$, $y$ isn’t a gateway vertex. Since $(a,y)$ is an edge, we must have $y\in\mathcal{S}$ or $y\in C$. In the former case since $y$ isn’t a gateway vertex, and in the latter case by Lemma 2.3, there exists $Q\in\mathcal{P}$ with $y\in Q$. Let $(a^{\prime},y)$ be the edge in $Q$ taking $y$ ”backward” along $Q$. We claim that $a^{\prime}\in A$. If $y\in\mathcal{S}$, then $a^{\prime}\in\mathcal{S}$ since $(a^{\prime},y)$ is an edge; if $y\in C$ then by Lemma 2.3, $a^{\prime}\in\mathcal{S}$ (the path $Q$ doesn’t contain two cut-vertices). Since $(a^{\prime},x^{\prime}=\Pi_{r}(a^{\prime}))$ lies in the cover graph since the $r$-projection of the $(a^{\prime},y)$ lies there, we get $x^{\prime}$ lies in the cover-graph, implying $a^{\prime}$ must lie in $A$. The above figure illustrates this. In short, there are $|Y|$ paths of $\mathcal{P}$ that contain vertices of $A$. Since $|Y|=|A|$ and since these paths are all vertex-disjoint, we conclude all vertices in $A$ are present in some path of $\mathcal{P}$. However, the gateway vertex $v^{\star}\in A$ and doesn’t lie on any path of $\mathcal{P}$. Contradiction. ∎
@@
  • $B=\{b~{}:~{}(x,b)\in P,~{}\text{for some}~{}x\in X,P\in\mathcal{P}\}$.
@@
As in the proof of Lemma 2.5, we have $X\cup Y\subseteq L_{k+1}$, and $B\subseteq L_{k+2}$ and that $A,X,B,Y$ all lie in the cover graph $G_{\mathcal{S},{\cal T}}$. As there, we crucially use the property that the projection of every $(A,Y)$ edge lies is an $(X,B)$ edge. Note that, we have $|A|=|X|$ and $|B|=|Y|$. However, unlike in Lemma 2.5, there can be two paths of $\mathcal{P}$ that may contain $x\in X$, and thus, $|B|$ may be larger in size than $|X|$. And this makes the proof slightly more complicated.
We now restrict our attention to these four sets. As a result, we only consider the cover graph $G_{A,B}$, which, recall, is the union of all paths from vertices in $A$ to vertices in $B$ (irrespective of whether they contain cut-elements). Both $A$ and $B$ lie in this graph by definition. We assert $X\subseteq G_{A,B}$. Take a vertex $x\in X$. If $x\in P$ for some path $P\in\mathcal{P}$ then it lies in $G_{A,B}$ by construction. We now show if $x$ is not any path in $\mathcal{P}$, we already have a contradiction. Indeed, the edge $(a,x)$ where $a=\Pi_{r}(x)$ isn’t in $F$ for otherwise, by Lemma 3.2 there would be a path containing $(a,x)$. Similarly $x\notin C$. So, $x\in\mathcal{S}$. Since $x\in L_{k+1}$ and $k<j-1$, there is at least one edge leaving it, and as before, this edge isn’t in $F$. So, $x$ is a gateway vertex which contradicts the choice of $k$. Thus, the subset $X$ lies in $G_{A,B}$. And since every $(A,Y)$ edge is an $r$-projection of an $(X,B)$ edge, we argue that $Y$ also lies in $G_{A,B}$.
@@ 4 Connections to monotonicity testing
The motivation for Conjecture 1.3 (and indeed, Theorem 1.1) is a deeper understanding of the problem of monotonicity testing of functions, a problem which, especially over the hypercube and hypergrid domains, has had a rich history of more than 25 years  [Ras99, EKK+00, GGL+00, DGL+99, LR01, FLN+02, HK03, AC06, HK08, ACCL07, Fis04, SS08, Bha08, BCSM12, FR10, BBM12, RRS+12, BGJ+12, CS13, CS14, CST14, BRY14a, BRY14b, CDST15, CDJS17, KMS18, BB21, CWX17, BCS18, BCS20, BKR20, HY22, BKKM23, BCS23b, BCS23a]. A function $f:\{0,1\}^{d}\to\{0,1\}$ is monotone if $\forall x,y\in\{0,1\}^{d}$ where $x\prec y$, $f(x)\leq f(y)$. The distance between two functions $f,g$ is |$|\{x:f(x)\neq g(x)\}|/2^{d}$, and the distance of $f$ to monotonicity, denote $\varepsilon_{f}$, is the minimum distance of $f$ to a monotone $g$. A function $f$ is said to be $\varepsilon$-far from monotone if $\varepsilon_{f}\geq\varepsilon$. The aim of a tester is to distinguish a monotone function from one that is ”far” from monotone. There is a special focus on non-adaptive monotonicity testers with one-sided error. These are testers that (i) always accept monotone functions, and (ii) make all their queries in advance. After a long line of work, this has been resolved (up to $\log d,\mathrm{poly}(\varepsilon^{-1})$ factors) to be $\Theta(\sqrt{d})$ [GGL+00, FLN+02, CS14, CST14, KMS18, CDST15, CWX17].
@@
In the next subsection, we give combinatorial interpretations to each of these statements. The reason for Conjecture 1.3 and a deeper study of hypercube routing was to get alternate proofs of Theorem 4.4. A big mystery of all these directed isoperimetric inequalities is the appearance of $\varepsilon_{f}$, the distance to monotonicity, as a ”directed version” of the variance of $f$. It appears as if $\varepsilon_{f}$ is the ”right measure” of directed volume. We hope that alternate proofs of Theorem 4.4 may shed some light on this mystery.